Section 2

Algebra

In case of several authors, Invited Speakers are marked with a ^{∗}.

Eric M. Friedlander: Geometry of Infinitesimal Group Schemes . . II  55

Sergei V. Ivanov: On the Burnside Problem for Groups of Even Exponent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 67

William M. Kantor: Simple Groups in Computational Group Theory . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 77

Gunter Malle: Spetses . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . II  87

Aleksandr V. Pukhlikov: Birational Automorphisms of Higher-Dimensional Algebraic Varieties . . . . . . . . . . . . . . . . . . . . . . . . . . . . II 97

Idun Reiten: Tilting Theory and Quasitilted Algebras . . . . . . . . . . . . . . II 109
Jeremy Rickard: The Abelian Defect Group Conjecture . . . . . . . . . . . II 121
Aner Shalev: Simple Groups, Permutation Groups, and Probability II 129

Documenta Mathematica · Extra Volume ICM 1998 · II ·

54

Documenta Mathematica · Extra Volume ICM 1998 · II ·